OpenStudy (anonymous):

A cafeteria serves three choices for lunch: turkey, chicken, or veggie. The turkey plate costs $4.80, the chicken plate costs $2.40, and the vegetable plate costs $1.20. If the lunches are mixed making 6 orders for $3.40 each, with twice as many chicken plates as veggie plates, find the number of each lunch plate mixed. I need to work on this with someone

OpenStudy (anonymous):

So we need variables for our unknowns. Let them be \(t, c, v\) for turkey, chicken, and vegetable respectively. We know \(2c=v\) since there are twice as many chicken plates as veggie plates. Can you think of any other equations we have?

OpenStudy (anonymous):

oops I meant \(c=2v\)

OpenStudy (anonymous):

yea so we do \[t+c+v=6\]\[2v+v+t=3.4(6)\]

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

\[t+v+2v=6\]\[4.8t+1.2(2v)=3.4(6)\]

OpenStudy (anonymous):

\(t+c+v=6\) is correct. \(c+v+t = 3.4(6)\) is incorrect. We need prices for all of them!

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

The second one should be: \(4.8t + 2.4c + 1.2v = 3.4(6)\) Ummm try not to substitute until you have all your equations or you'll start confusing me, lol

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

and then we substitute for the c's right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

k thanks wio! :D

OpenStudy (anonymous):

You can always check your solutions here: http://www.wolframalpha.com/input/?i=c%3D2v%2C+t%2Bc%2Bv%3D6%2C+4.8t%2B2.4c%2B1.2v%3D3.4%286%29